# The Sunburst Arc with JWST. IV. The importance of interaction, turbulence, and feedback for Lyman-continuum escape

### v2

@@ 3.1 Correction of Uncertainties
We estimated the statistical uncertainty in the in every wavelength bin in the Level 3 datacubes as follows. We first manually constructed masks covering regions of empty sky. These masks were placed to minimize line emission from the source galaxy, but were also kept well clear of edges of each pointing, since these regions are noisier and not representative of the noise properties of the spaxels covering the target. The masks are shown in Figure 13.
@@ 3.2 Stacked spectrum of the LyC leaking knot
Of the twelve gravitationally lensed images of the LCE cluster, five are captured by the NIRSpec IFU pointings. Since these are all physically the same object, we combined their spectra by extracting each in a pixels aperture, normalized each spectrum by its median value to standardize their differently magnified fluxes, before stacking them by taking a flux-weighted average. All quantities derived from this spectrum are thus given in relative or units.
Figure 3 shows the full stacked, rest-frame blue and optical spectrum of the LCE cluster. The spectrum contains a large number of emission lines; we have identified 59 nebular lines, of which 30 are Hydrogen recombination lines (including about 20 Balmer lines and 10 Paschen lines). The spectra also feature a set of broad, blended, stellar Wolf-Rayet line complexes that were discussed in T. E. Rivera-Thorsen et al. (2024). The Figures in Appendix B identify all these features. To enable conversion to physical units, we report the measured and the demagnifed flux of H for the image of the LCE that appears in image 4 of the lensed galaxy: the measured flux is , and the demagnfied flux is , assuming the calculated magnfication of from K. Sharon et al. (2022).
@@
For spectral line modeling and generation of maps of line flux and kinematics, we used the Julia package CubeFitter.jl (T. E. Rivera-Thorsen, 2025)aaahttps://github.com/thriveth/CubeFitter.jl, a software package that quickly and flexibly models emission lines in spectroscopic datacubes. CubeFitter is written in the Julia programming language (J. Bezanson et al., 2017), to strike a balance between computation speed and simple, flexible development. Julia has proven for our tasks to be faster than Python, while taking up a number of lines of code and a level of coding complexity similar to Python code of equal functionality. CubeFitter relies on a number of existing Julia packages for its core functionality, most importantly the general curve fitting package GModelFit.jlbbbhttps://gcalderone.github.io/GModelFit.jl for model building infrastructure and fitting functionality, Measurements.jlccchttps://juliaphysics.github.io/Measurements.jl for propagation of uncertainties, and optionally VoronoiBinning.jldddhttps://github.com/Michael-Reefe/VoronoiBinning.jl, a Julia implementation of the original Vorbin package (M. Cappellari & Y. Copin, 2003).
The CubeFitter.jl package loads a FITS datacube into a data structure with relevant, partially user-provided, metadata such as instrument resolving power, data units, and more. The program ships with a list of emission lines, mainly taken from Drew Chojnowski’s online line listeeehttp://astronomy.nmsu.edu/drewski/tableofemissionlines.html, in turn compiled mainly from data from NIST (A. Kramida et al., 2013). A few additional lines are taken from other places, see the package documentation. Alternatively, the user can pass their own line list to the code. The package incorporates the tables of spectral resolution for NIRSpecfffavailable for download at https://jwst-docs.stsci.edu/jwst-near-infrared-spectrograph/nirspec-instrumentation/nirspec-dispersers-and-filters#gsc.tab=0. These are pre-launch estimates; we are not aware of in-orbit measurements.
@@
Data regions with undetected lines contribute no signal to the fit, increase computation time, and unnecessarily increase uncertainties of the resulting fit parameters. To avoid this, the code performs a quick, numerical estimate of the signal-to-noise ratio of each line in each spaxel, and discards any lines below a (settable) S/N threshold, inserting a NaN value in the resulting flux maps for the given line.
The code has rudimentary support for modeling two kinematic components in each line, but at the time of writing this, it is not yet possible to combine this with the use a custom selection of lines to compute the kinematics.
@@ 3.4.1 Line fitting strategy
We picked a number of the strongest lines on which to base the kinematics. In the case of F100LP/G140H, we used [O II], H, and [O III]; and for F170LP/G235H we used [O III], H, and [N II] (blended features must be fitted together to get correct fluxes). These were then used to build one single model, which was fit to the observed data. We have not attempted to account for the possible presence of multiple gas components in each line, and we have not attempted to account for possible kinematic differences between various gas phases. A more detailed and in-depth kinematic modeling of the gas is deferred to a follow-up paper.
@@ 3.4.2 Masking
Since the kinematics in most pixels is dominated by the [O III] 4960,5008 Å doublet, we produced a stacked [O III] S/N map by assuming a fixed line ratio of and, once this fixed line ratio was corrected for, taking the mean of the four line maps (each doublet member in each of the two settings). Uncertainties were propagated in the standard way. We then created a S/N map, and excluded all spaxels with for this stacked [O III] map.
@@ 3.4.3 Combined maps
3.5 Determining the dust attenuation and ionization properties
We have accounted for dust attenuation in the galaxy using a standard starburst attenuation law (D. Calzetti et al., 2000). We have computed using the Balmer decrement of H/H following standard methods as outlined in e.g. A. Domínguez et al. (2013), and using intrinsic line ratios from D. E. Osterbrock & G. J. Ferland (2006) assuming and . In the very high density region surrounding the “Godzilla” region (see S. Choe et al., 2025), the emission from partially ionized regions (L. Dai, 2024) or broad H emission from the power-law winds of the extremely lensed Godzilla object may lead to an elevated H/H line ratio and thus an overestimation of ; a more thorough treatment of dust attenuation in this region was done by S. Choe et al. (2025), who found a strongly elevated from H in this clump, while a selection of other Balmer and Paschen lines unaffected by this effect yielded a value of only half that. We therefore advise caution when interpreting the values found from H near the Godzilla region. However, the rest of the areas covered by the IFU are more typical for warm starburst ISM, and thus a H based determination of is appropriate for this region.
@@ 3.7.1 Masking and de-magnification map
To be able to imitate physical properties of the galaxy as they would have been observed in a slit spectrum without gravitational lensing, we extracted the global spectroscopic properties of the galaxy by the following method. First, we constructed a mask which contained only the one lensed image that covers the largest part of the source galaxy. The region, which is found in P1, is shown in Figure 4. This galaxy image is bounded in both directions along the arc by crossing critical curves, and in the off-arc direction predominantly by the [O III] S/N based mask from subsubsection 3.4.2.
The location of the critical curves is derived from the lensing model of K. Sharon et al. (2022). To compensate for differential magnification within the mask, we used magnification maps derived from the same work. We used the Python package Reproject to project the magnification maps onto the WCS of the drizzled NIRSpec cube of P1. We then applied the combined S/N and critical curve mask to the magnification map as well. The inverse of the resulting masked map can then be used to compensate for differential magnification. We note that the exact delineation of the mask along the critical curve is of limited importance: Magnification is high in the vicinity of the critical curve, and since we weigh the data by the inverse magnification, a few pixels too many or too few in this region will have very limited impact on the resulting values of integrated and weighted quantities.
A more thorough treatment of a magnification-weighted, integrated spectrum and comparisons to non-resolved starburst galaxies at similar redshifts is deferred to future work. Here, we have used it to derive global values of properties derived per-spaxel in this work; including redshift, ionization parameter, [S II] deficiency (see subsubsection 3.7.3), and more. An important caveat to these properties is that, as is evident from Figure 2, the image shown in Figure 4, while covering most of the galaxy, is not a complete image. All global properties are given under the assumption that the included regions fairly represent the statistical properties of the galaxy.
@@ 3.7.2 Global kinematic properties
A common means to characterize the kinematical properties of a galaxy is by comparing the shearing velocity to the intrinsic velocity dispersion (K. Glazebrook, 2013; E. C. Herenz et al., 2016; A. Bik et al., 2022; E. C. Herenz et al., 2025). is defined as the difference in centroid velocity between the spaxels with the most extreme values. is defined as the flux-weighted velocity dispersion averaged over all spaxels. Following E. C. Herenz et al. (2016), but substituting [O III]5008 for H, these are written as:
Following E. C. Herenz et al., we have used the 5th and 95th percentiles of velocities to avoid outliers from skewing the results, while still sampling the extreme ends of the velocity distribution. We opted to base on [O III]5008 rather than H, because in our dataset, the former both intrinsically has higher S/N than the latter, and because this line at the redshift of the Sunburst Arc falls in the overlapping range between G140H and G235H and thus was observed in both, effectively doubling the exposure time. When computing , we have not made any attempt to correct for beam smearing such as has been done in the ground based studies of e.g. K. Glazebrook (2013); E. C. Herenz et al. (2016); A. Bik et al. (2022); E. C. Herenz et al. (2025). The effective spatial resolution of JWST/NIRSpec for this lensed target is so high that the problem is negligible. Additionally, the beam smearing correction in these works was motivated by a central spike in line width in those kinematics maps caused by the atmospheric blurring, leading to an overestimation of the brightness-weighted velocity dispersion in these works. We do not observe such an effect, and consequently do not expect the same skewing of .
@@ 3.7.3 Ionization properties and diagnostic line ratios
Using the single-galaxy mask and inverse magnification weight map, we computed diagnostic line ratios and derived properties that are often used in the literature, so that the Sunburst Arc can be directly compared to unlensed galaxies at similar redshifts. These also provided the opportunity to compare the properties derived from the stacked spectrum of the LCE, to a spectrum as it would more realistically look if unlensed and unresolved. This could tell us whether the Sunburst Arc is globally similar to the LCE galaxies found in LCE surveys (K. Nakajima et al., 2020; J. Cooke et al., 2014; T. J. Fletcher et al., 2019; Y. Liu et al., 2023; B. Wang et al., 2021; S. R. Flury et al., 2022b; X. Xu et al., 2023; S. R. Flury et al., 2022a).
We calculated each of these diagnostics by taking the inverse-magnification weighted sum of each emission line within the single-image mask discussed in Sect. 3.7.1. Flux maps were sigma clipped. We then used these weighted, masked and integrated line fluxes to compute derived properties. We picked a superset of the spatially mapped diagnostics discussed in Sect. 3.7.3. In addition to those, we computed the sulphur deficiency as the difference between and the locus described in B. Wang et al. (2021).
In a similar manner, we computed global values of and derived properties.
@@ 4.1 Line flux maps
Figure 5 and Figure 6 show the emission intensity maps for the majority of the lines used in the diagnostic maps that follow. Each line map is colored to loosely correspond to the rest-frame wavelength of the line. The masks described in subsubsection 3.4.2 are applied here and in all later maps, unless otherwise stated.
@@ 4.2 Kinematics
The right column of Figure 7 shows the velocity-space FWHM. Of particular interest is the dramatic line broadening immediately to one side of the multiple images of the LCE but, interestingly, not actually coincident with it. Comparing to the middle column of Figure 7, we see that this region of high FWHM is coincident with a slight spike in outflow velocity. This offset kinematic spike might very well originate in a separate outflowing gas component which is automatically incorporated into our single-component model. In P2+3, this region appears particularly dramatic and elongated immediately NE of the LCE image 1.8 (in the terminology of K. Sharon et al., 2022), immediately next to the object “Godzilla”. This region is subject to extreme magnification and shearing due to a low surface brightness lensing foreground galaxy S. Choe et al. (2025), but despite this extreme magnification and distortion, it is still another lensed image of the same region as in P1.
@@ 4.3 Dust and ionization
In the center-left column, we show maps of derived as described in subsection 3.5 from the H/H ratio and a Calzetti attenuation law. The stripey patterns showing in these panels are an artifact of the undersampled PSF. We see that the dust attenuation is quite uniform across the arc, except in the “Godzilla” feature marked in Figure 1, and its much fainter counterimages in both P1 and the lower left of P2+3 (see S. Choe et al., 2025, for a further discussion).
@@
Finally, the right panels show another ionization diagnostic, , which has been proposed as an improved tracer of LyC escape due to the lower ionization potential of [S II] compared to [O II], meaning that these lines more accurately trace the presence or absence of neutral gas, where O+ ions also are present in substantial amounts inside H II regions and the photodissociation regions at their boundaries. While the [S II] line may be fainter than [O II], the diagnostic is based entirely on lines very close in wavelength and thus independent of dust model. For the same reason, it does not display the stripey artifacts of the other maps in this figure. It is to our knowledge the first time it has been possible to map both log [S II]/H and LyC emission in any galaxy and thus test its strength as a tracer on a local scale. We note that there is a strong similarity between the spatial distribution of high- and low-ionization regions by the two methods.
@@ 4.4 Nitrogen spike
Figure 9 shows a maps of the diagnostic N2. This diagnostic is powerful because of its short wavelength baseline, eliminating dust model dependence; on the other hand, it can only show the N/H abundance, not the relative abundance of N relative to other elements. We note the conspicuous spikes in N/H apparent in these maps at locations next to, but not coincident with, the LCE cluster.
@@ 5.1 Kinematics
If we adopt the median inclination angle reached above, value, this gives a correction factor to of , yielding a corrected value of . In (N. M. Förster Schreiber & S. Wuyts, 2020; A. Bik et al., 2022), a value of is adopted as the demarkation above which they consider a galaxy to be rotation dominated. This value is, for the Sunburst Arc galaxy, reached at angles of (see section 4). In a random distribution, one finds about 30% of galaxies in this inclination angle range, meaning that the galaxy, assuming no knowledge of its orientation, has about a 70% probability of being dispersion rather than rotation dominated, and for the majority of the remaining 30% not very strongly rotation dominated.
This paints a picture of a galaxy which is rotating, but also undergoing significant turbulence. From the FWHM maps in Figure 7, the velocity disperson does not spatially correlate with strong star formation activity, suggesting that this strong dispersion may be due to tidal disruptions by the Northern companion seen in Figure 2.

### v3

@@ 3.1 Correction of Uncertainties
We estimated the statistical uncertainty in every wavelength bin in the Level 3 datacubes as follows. We first manually constructed masks covering regions of empty sky. These masks were placed to minimize line emission from the source galaxy, but were also kept well clear of edges of each pointing, since these regions are noisier and not representative of the noise properties of the spaxels covering the target. The masks are shown in Figure 13.
@@ 3.2 Stacked spectrum of the LyC leaking knot
Of the twelve gravitationally lensed images of the LCE cluster, five are fully captured by the NIRSpec IFU pointings. Since these are all physically the same object, we combined their spectra by extracting each in a pixels aperture, normalized each spectrum by its median value to standardize their differently magnified fluxes, before stacking them by taking a flux-weighted average. All quantities derived from this spectrum are thus given in relative or units.
Figure 3 shows the full stacked, rest-frame blue and optical spectrum of the LCE cluster. The spectrum contains a large number of emission lines; we have identified 59 nebular lines, of which 30 are Hydrogen recombination lines (including about 20 Balmer lines and 10 Paschen lines). The spectra also feature a set of broad, blended, stellar Wolf-Rayet line complexes that were discussed in riverathorsen2024. The Figures in Appendix B identify all these features. To enable conversion to physical units, we report the measured and the demagnifed flux of H for the image of the LCE that appears in image 4 of the lensed galaxy: the measured flux is , and the demagnfied flux is , assuming the calculated magnfication of from K. Sharon et al. (2022).
@@
For spectral line modeling and generation of maps of line flux and kinematics, we used the Julia package CubeFitter.jl (riverathorsen2025cubefitter)111https://github.com/thriveth/CubeFitter.jl, a software package that quickly and flexibly models emission lines in spectroscopic datacubes. CubeFitter is written in the Julia programming language (bezanson2017julia), to strike a balance between computation speed and simple, flexible development. Julia has proven for our tasks to be faster than Python, while taking up a number of lines of code and a level of coding complexity similar to Python code of equal functionality. CubeFitter relies on a number of existing Julia packages for its core functionality, most importantly the general curve fitting package GModelFit.jl222https://gcalderone.github.io/GModelFit.jl for model building infrastructure and fitting functionality, Measurements.jl333https://juliaphysics.github.io/Measurements.jl for propagation of uncertainties, and optionally VoronoiBinning.jl444https://github.com/Michael-Reefe/VoronoiBinning.jl, a Julia implementation of the original Vorbin package (capellari2003).
The CubeFitter.jl package loads a FITS datacube into a data structure with relevant, partially user-provided, metadata such as instrument resolving power, data units, and more. The program ships with a list of emission lines, mainly taken from Drew Chojnowski’s online line list555http://astronomy.nmsu.edu/drewski/tableofemissionlines.html, in turn compiled mainly from data from NIST (nist_asd). A few additional lines are taken from other places, see the package documentation. Alternatively, the user can pass their own line list to the code. The package incorporates the tables of spectral resolution for NIRSpec666available for download at https://jwst-docs.stsci.edu/jwst-near-infrared-spectrograph/nirspec-instrumentation/nirspec-dispersers-and-filters#gsc.tab=0. These are pre-launch estimates; although in-orbit measurements have recently been made (shajib2025arxiv), they have not yet yielded data products suitable for implementation in code like this..
@@
Data regions with undetected lines contribute no signal to the fit, increase computation time, and unnecessarily increase uncertainties of the resulting fit parameters. To avoid this, the code performs a quick, numerical estimate of the signal-to-noise ratio of each line in each spaxel, and discards any lines below a (settable) S/N threshold, inserting a NaN value in the resulting flux maps for the given line. The code has rudimentary support for modeling two kinematic components in each line, but at the time of writing this, it is not yet possible to combine this with the use of a custom selection of lines to compute the kinematics.
@@ 3.4.1 Line fitting strategy
We picked a number of the strongest lines on which to base the kinematics. In the case of F100LP/G140H, we used [O II], H, and [O III]; and for F170LP/G235H we used [O III], H, and [N II] (blended features must be fitted together to get correct fluxes). These were then used to build one single model, which was fit to the observed data. We have not attempted to account for the possible presence of multiple gas components in each line, and we have not attempted to account for possible kinematic differences between various ionization zones. A more detailed and in-depth kinematic modeling of the gas is deferred to a follow-up paper.
@@ 3.4.2 Masking
Since the kinematics in most pixels is dominated by the [O III] 4960,5008 Å doublet, we produced a stacked [O III] S/N map by assuming a fixed line ratio of (vanhoof2018) and, once this fixed line ratio was corrected for, taking the mean of the four line maps (each doublet member in each of the two settings). Uncertainties were propagated in the standard way. We then created a S/N map, and excluded all spaxels with for this stacked [O III] map.
@@ 3.4.3 Combined maps
3.5 Dust attenuation maps
We have accounted for dust attenuation in the galaxy using a standard starburst attenuation law (calzetti2000). We have computed using the Balmer decrement of H/H following standard methods as outlined in e.g. dominguez2013, and using intrinsic line ratios from osterbrock assuming and . In the very high density region surrounding the “Godzilla” region (see choe2025), the emission from partially ionized regions (dai2024arxiv) or broad H emission from the power-law winds of the extremely lensed Godzilla object may lead to an elevated H/H line ratio and thus an overestimation of ; a more thorough treatment of dust attenuation in this region was done by choe2025, who found a strongly elevated from H in this clump, while a selection of other Balmer and Paschen lines unaffected by this effect yielded a value of only half that. We therefore advise caution when interpreting the values found from H near the Godzilla region. However, the rest of the areas covered by the IFU are more typical for warm starburst ISM, and thus a H based determination of is appropriate for this region.
@@ 3.7.1 Masking and de-magnification map
To emulate properties of the galaxy as they would have been observed in a slit spectrum without gravitational lensing, we extracted the global spectroscopic properties of the galaxy by the following method. First, we constructed a mask which contained only the one lensed image that covers the largest part of the source galaxy. The region, which is found in P1, is shown in Figure 4. This galaxy image is bounded in both directions along the arc by crossing critical curves, and in the off-arc direction predominantly by the [O III] S/N based mask from subsubsection 3.4.2.
The location of the critical curves is derived from the lensing model of sharon2022. To compensate for differential magnification within the mask, we used magnification maps derived from the same work. We used the Python package Reproject to project the magnification maps onto the WCS of the drizzled NIRSpec cube of P1. We then applied the combined S/N and critical curve mask to the magnification map as well. The inverse of the resulting masked map can then be used to compensate for differential magnification. We note that the exact delineation of the mask along the critical curve is of limited importance: Magnification is high in the vicinity of the critical curve, and since we weigh the data by the inverse magnification, a few pixels too many or too few in this region will have very limited impact on the resulting values of integrated and weighted quantities.
A more thorough treatment of a magnification-weighted, integrated spectrum and comparisons to non-resolved starburst galaxies at similar redshifts is deferred to future work. Here, we have used it to derive global values of properties derived per-spaxel in this work; including redshift, ionization parameter, [S II] deficiency (see subsubsection 3.7.3), and more. An important caveat to these properties is that the regions included in the mask in pointing 1 are not a fully complete image of the galaxy. As is evident from Figure 2, the image shown in Figure 4 covers most of the galaxy, but not all. All global properties are given under the assumption that the included regions are an accurate representation of the global properties of the galaxy.
@@ 3.7.2 Global kinematic properties
A common means to characterize the kinematical properties of a galaxy is by comparing the shearing velocity to the intrinsic velocity dispersion (glazebrook2013; herenz2016; bik2022; herenz2025). is defined as the difference in centroid velocity between the spaxels with the most extreme values. is defined as the flux-weighted velocity dispersion averaged over all spaxels. Following herenz2016, but substituting [O III]5008 for H, these are written as:
Following herenz2016, we have used the 5th and 95th percentiles of velocities to avoid outliers from skewing the results, while still sampling the extreme ends of the velocity distribution. We opted to base on [O III]5008 rather than H, because in our dataset, the former both intrinsically has higher S/N than the latter, and because this line at the redshift of the Sunburst Arc falls in the overlapping range between G140H and G235H and thus was observed in both, effectively doubling the exposure time. When computing , we have not made any attempt to correct for beam smearing such as has been done in the ground based studies of e.g. glazebrook2013; herenz2016; bik2022; herenz2025. The effective spatial resolution of JWST/NIRSpec for this lensed target is so high that the problem is negligible. Additionally, the beam smearing correction in these works was motivated by a central spike in line width in those kinematics maps caused by the atmospheric blurring, leading to an overestimation of the brightness-weighted velocity dispersion in these works. We do not observe such an effect, and consequently do not expect the same skewing of .
@@ 3.7.3 Ionization properties and diagnostic line ratios
Using the single-galaxy mask and inverse magnification weight map, we have computed global values of diagnostic line ratios and derived properties that are often used in the literature, so that the Sunburst Arc can be directly compared to observations of unlensed galaxies at similar redshifts. These also provided the opportunity to compare the properties derived from the stacked spectrum of the LCE, to a spectrum as it would more realistically look if unlensed and unresolved. This could tell us whether the Sunburst Arc is globally similar to the LCE galaxies found in LCE surveys (e.g., nakajima2020; cooke2014; fletcher2018; liu2023; wang2021; xu2023; flury2022a; flury2022b).
We calculated each of these diagnostics by extracting a inverse magnification weighted spectrum in the single-image mask discussed in Sect. 3.7.1. We then used these weighted, masked and integrated line fluxes to compute derived properties. In addition to the global kinematic properties and the spatially mapped diagnostics discussed in Sect. 3.7.3, we computed the sulphur deficiency as the difference between and the locus described in wang2021. In a similar manner, we computed global values of and derived properties.
@@ 4.1 Line flux maps
Figure 5 and Figure 6 show the emission intensity maps for the majority of the lines used in the diagnostic maps that follow. Each line map is colored to loosely correspond to the rest-frame wavelength of the line. The masks described in subsubsection 3.4.2 are applied here and in all subsequent maps, unless otherwise stated.
@@ 4.2 Kinematics
The right column of Figure 7 shows the velocity-space FWHM. Of particular interest is the dramatic line broadening immediately to one side of the multiple images of the LCE but, interestingly, not actually coincident with it. Comparing to the middle column of Figure 7, we see that this region of high FWHM is coincident with a slight spike in outflow velocity. This offset kinematic spike might very well originate in a separate outflowing gas component which is automatically incorporated into our single-component model. In P2+3, this region appears particularly dramatic and elongated immediately NE of the LCE image 1.8 (in the terminology of sharon2022), immediately next to the object “Godzilla”. This region is subject to extreme magnification and shearing due to a low surface brightness lensing foreground galaxy (choe2025), but despite this extreme magnification and distortion, it is still another lensed image of the same region as in P1.
@@ 4.3 Dust and ionization
In the center-left column, we show maps of derived as described in subsection 3.5 from the H/H ratio and a Calzetti attenuation law. The stripey patterns showing in these panels are an artifact of the undersampled PSF. We see that the dust attenuation is quite uniform across the arc, except in the “Godzilla” feature marked in Figure 1, and its much fainter counterimages in both P1 and the lower left of P2+3 (see choe2025, for a further discussion).
@@
Finally, the right panels show another ionization diagnostic, , which has been proposed as an improved tracer of LyC escape due to the lower ionization potential of [S II] compared to [O II], meaning that these lines more accurately trace the presence or absence of neutral gas, where O+ ions also are present in substantial amounts inside H II regions and the photodissociation regions at their boundaries. While the [S II] line may be fainter than [O II], the diagnostic is based entirely on lines very close in wavelength and thus independent of dust model. For the same reason, it these stripey artifacts are suppressed in this map, although still visible. It is to our knowledge the first time it has been possible to map both log [S II]/H and LyC emission in any galaxy and thus test its strength as a tracer on a local scale. We note that there is a strong similarity between the spatial distribution of high- and low-ionization regions by the two methods.
@@ 4.4 Nitrogen spike
Figure 9 shows a maps of N2 = [N II]/H. We note the conspicuous regions of elevated [N II]/H at locations next to, but not coincident with, the LCE cluster. In these regions, the elevated [N II] might be due to elevated N/H, or due to collisional excitation from shocks. Spatially, they closely match the regions of larger line FWHM and a slight rise in blueshift as seen in Figure 7; something which could support the collisional interpretation. The region is not coincident with any star clusters we have been able to identify in our NIRCam images or archival HST images; we speculate that it might be the result of an anisotropic outflow from the LCE cluster.
@@ 5.1 Kinematics
If we adopt the median inclination angle reached above, value, this gives a correction factor to of , yielding a corrected value of . In (schreiber2020; bik2022), a value of is adopted as the demarkation above which they consider a galaxy to be rotation dominated. This value is, for the Sunburst Arc galaxy, reached at angles of (see section 4). In a random distribution, one finds about 30% of galaxies in this inclination angle range, meaning that the galaxy, assuming no knowledge of its orientation, has about a 70% probability of being dispersion rather than rotation dominated, and for the majority of the remaining 30% not very strongly rotation dominated.
This paints a picture of a galaxy which is rotating, but also undergoing significant turbulence. From the FWHM maps in Figure 7, the velocity disperson does not spatially correlate with strong star formation activity (apart from localized and directional outflows near the LCE, cf. section 4.2 and 4.4), suggesting that this strong dispersion, rather than stemming from star formation feedback, may be due to outside influences such as tidal disruptions by the Northern companion seen in Figure 2.